OpenStudy (anonymous):

How would I solve the system of equation for: x + 4y + z = 12 y - 3z = -7 z = 3

OpenStudy (anonymous):

x+4y+3=12 y-9=-7 So y=2 x=1 x=3

OpenStudy (anonymous):

Can I have a metal

OpenStudy (anonymous):

Substitute the value of z in the second euation to get y,thn substitute both these in first to get x y -3(3)=-7. x y-9=-7 y=2.